OpenStudy (kaptain_mittens):

Please Help! Medal and Fan! Simplify square root of 2 over cube root of 2

OpenStudy (anonymous):

2*2 is squared

OpenStudy (anonymous):

and 2*2*2 is cubed

OpenStudy (kaptain_mittens):

yeah i know that but this is square root 2 over 3^square root 2

OpenStudy (kaptain_mittens):

@texaschic101

OpenStudy (kaptain_mittens):

@Directrix

OpenStudy (kaptain_mittens):

@dan815

OpenStudy (kaptain_mittens):

@ganeshie8

jimthompson5910 (jim_thompson5910):

hint: \[\Large \frac{\sqrt{2}}{\sqrt[3]{2}} = \frac{2^{1/2}}{2^{1/3}}\]

OpenStudy (kaptain_mittens):

ok so make those common denominators, right?

jimthompson5910 (jim_thompson5910):

you subtract the exponents for the next step

jimthompson5910 (jim_thompson5910):

Rule: \[\LARGE \frac{x^y}{x^z} = x^{y-z}\]

OpenStudy (kaptain_mittens):

2^1/6

jimthompson5910 (jim_thompson5910):

yes, \[\Large \frac{\sqrt{2}}{\sqrt[3]{2}} = \frac{2^{1/2}}{2^{1/3}}\] \[\Large \frac{\sqrt{2}}{\sqrt[3]{2}} = 2^{1/2 - 1/3}\] \[\Large \frac{\sqrt{2}}{\sqrt[3]{2}} = 2^{3/6 - 2/6}\] \[\Large \frac{\sqrt{2}}{\sqrt[3]{2}} = 2^{1/6}\] \[\Large \frac{\sqrt{2}}{\sqrt[3]{2}} = \sqrt[6]{2}\]

OpenStudy (kaptain_mittens):

Thanks for coming through for me again man, much appreciated

jimthompson5910 (jim_thompson5910):

you're welcome